OpenStudy (anonymous):

Given the following endpoint (A) and midpoint (M), find the other endpoint (B). M(2, 3) A(0, -1) 1) (4, 7) 2) (-1, 0) 3) (3, 2) 4) (7, 4)

OpenStudy (anonymous):

(4, 7)

OpenStudy (anonymous):

the formula is B=2M-A B=2(2,3)-(0,-1) =(4,6)-(0,-1) =(4,7)